# LEASE: Offline Preference-based Reinforcement Learning with High Sample Efficiency

### v1

@@ V-A Uncertainty-aware Pseudo-label Selection
where $\kappa_{p}$ and $\kappa_{\tau}$ are thresholds of confidence and uncertainty respectively. $f(\sigma_{0}^{u},\sigma_{1}^{u})=1$ denotes the generated data $(\sigma_{0}^{u},\sigma_{1}^{u})$ is selected, and $f(\sigma_{0}^{u},\sigma_{1}^{u})=0$ denotes the data is not selected. Then combing Eq. (6), the reward model can be optimized through minimizing
^$\widehat{\mathcal{L}}_{R}(\psi)=\frac{1}{\widetilde{N}_{u}}\sum\nolimits_{j=1}^{{N}_{u}}f\left(\sigma_{0}^{u},\sigma_{1}^{u}\right)^{(j)}L\left((\sigma_{0}^{u},\sigma_{1}^{u})^{(j)},\widehat{y}^{j}\right)$ (9)
@@
where $\widetilde{N}_{u}$ is the number of generated dataset after screening, and ^$\widehat{\mathcal{L}}_{R}(\psi)$ is the empirical risk. It contains two parts: labeled loss ^$\widehat{\mathcal{L}}_{l}(\psi)$ and unlabeled loss ^$\widehat{\mathcal{L}}_{u}(\psi)$. Note that the label in unlabeled loss is pseudo-label, which may be different from the true label. Therefore, there is the gap between the unlabeled loss with pseudo-label ^$\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label ^$\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$. Before bounding this gap, we firstly give the below assumption without loss of generality.
@@
|^$\big{|}\widehat{\mathcal{L}}_{u}(\psi)-\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\big{|}\leq\eta\Omega.$ (10)
The proof of Proposition 1 can be found in Appendix A-A. This gap is mainly influenced by the term $\eta$. Through selecting high confidence and low uncertainty samples, the $\eta$ can be significantly reduced. Combining Eqs. (5) and (8), we find that if $p(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})$ is small, the $L((\sigma_{0}^{u},\sigma_{1}^{u}),\widehat{y})$ would become larger. Too large error can lead to unstable training and low performance. Therefore, through selecting mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$, a more accurate subset of pseudo-labels can be used in reward training, and through setting a higher confidence threshold for pseudo labeling, a lower prediction error and the stability of training can be guaranteed.
V-B Generalization Bound for Reward Model
Before developing generalization bound, we define the expected error $\mathcal{L}_{R}(\psi)$ with respect to the reward model $R(s,a;\psi)$: $\mathcal{L}_{R}(\psi)=\mathbb{E}_{(\sigma_{0},\sigma_{1},y)\sim\mathcal{D}}[L((\sigma_{0},\sigma_{1}),y)]$, where $\mathcal{D}$ is data distribution. The reward model is trained through minimizing the empirical error ^$\widehat{\mathcal{L}}_{R}(\psi)$ in Eq. (9). Through developing generalization error bound, we can analyse the factors that influence generalization ability.
Similar to the previous generalization bound works [29], Rademacher complexity, which measures the richness of a certain hypothesis space [19], is introduced firstly. The definition is given below:
@@ V-B Generalization Bound for Reward Model
^$\widehat{\Re}_{\widehat{\mathcal{S}}}(\mathcal{F})=\mathbb{E}_{\bm{\sigma}}\big{[}\sup_{f\in\mathcal{F}}\frac{1}{N_{s}}\sum\nolimits_{i=1}^{N_{s}}\sigma_{i}f\left(z_{i}\right)\big{]},$ (11)
@@
In general, lower Rademacher complexity corresponds to better generalization performance. Then, based on the above definition and Proposition 1, we develop the new generalization bound for reward model trained through Eq. (9). The specific theorem is given below.
@@
The proof of Theorem 1 can be found in Appendix A-B. Theorem 1 indicates that the expected error $\mathcal{L}_{R}(\psi)$ is bounded by empirical error ^$\widehat{\mathcal{L}}_{R}(\psi)$, pseudo-labeling error $\eta\Omega$, Rademacher complexity and constant terms. The constant term can become small when the number of unlabeled data increases, but it may cause unstable training when labels are inaccurate. According to Proposition 1, empirical error and pseudo-labeling error can be reduced through selecting mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$, thus the better generalization ability, that is tighter upper bound, can be achieved. Note that Theorem 1 has generality and can be applicable to methods that train reward model using pseudo-labels. The difference among various methods may lie in how they train reward model to improve the accuracy of pseudo-labels, that is reducing $\eta$. If one method fails to reduce the pseudo-labeling error $\eta$, the upper generalization bound would be looser than that of no data augmentation.
VI Policy Learning
This paper directly uses the proposed offline RL algorithms, such as CQL [4] and IQL [25], to perform policy learning. To solve the problem of high cost for labeling preference, we propose a offLine prEference-bAsed RL with high Sample Efficiency algorithm (LEASE). This section aims to describe the details of LEASE and establish the theory for safe policy improvement guarantee.
VI-A The Implementation Details of LEASE
@@
5:     if $t$ rollout frequency $=0$ then
6:        Generate synthetic $H$-step rollouts by ${T}_{\varphi}$. Add transition trajectory into $\mathcal{D}_{u}$.
@@
8:     if reward update condition is TRUE then
9:        Give pseudo label for unlabeled dataset $\mathcal{D}_{u}$ based on pre-trained reward model.
10:        Select high confidence and low uncertainty unlabeled data according to Eq. (8).
11:        Train the reward model based on $\mathcal{D}_{l}$ and $\mathcal{D}_{u}$ according to Eq. (9).
12:     end if
13:     Train policy using offline RL algorithms, such as CQL and IQL.
14:  end for
@@
Most offline RL algorithms are based on Eq. (1). They proposed various algorithms to solve the problem of distribution shift. In offline PbRL setting, the reward is unknown for agent. The accuracy of reward model directly influence the performance of policy. Moreover, the unlabeled data are generated through the current policy interaction with learned transition model. Therefore, the reward model and policy influence each other. In practical implementation, the policy, $Q$-function, reward and transition model are parameterized by $\pi_{\theta}$, $Q_{\omega}$, $R_{\psi}$ and $T_{\varphi}$ respectively.
Algorithm 1 gives pseudocode for LEASE. The goal of LEASE is to learn better reward model $R_{\psi}$ from fewer labeled preference dataset $\mathcal{D}_{l}$ and learn better policy $\pi_{\theta}$ from offline dataset $\mathcal{D}_{\text{offline}}$ without interaction with environment. The key of LEASE is how to train reward model in policy learning process. Notably, the reward model only update once instead of updating constantly in this process. The reward update condition is set to when the number of unlabeled data in buffer reaches the maximum buffer capacity. The update time is influenced by rollout frequency and rollout batch size. The reward update condition, rollout length $H$ and selecting mechanism $f(\sigma_{1},\sigma_{2})$ are the important factors for reward model.
VI-B Safe Policy Improvement Guarantee
Offline RL aims to guarantee $\xi$-safe policy improvement over the behavior policy $\mu$ (the policy used to collect offline dataset), i.e. $J(\widehat{\pi},R^{*})\geq J(\mu,R^{*})-\xi$ [4]. This part develops theoretical guarantee of policy improvement for LEASE, that is giving the bound for $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$, and further analyzes sample complexity when reward model is unknown. Firstly, we propose a new single-policy concentrability coefficient for PbRL based on state-action pair instead of trajectory like [18].
@@ VI-B Safe Policy Improvement Guarantee
$\mathscr{C}_{\mathcal{R}}(\pi)=\sup_{\widehat{R}\in\mathcal{R}}\Bigg{|}\frac{\mathbb{E}_{(s,a)\sim d_{T}^{\pi}}\big{[}R^{*}(s,a)-\widehat{R}(s,a)\big{]}}{\mathbb{E}_{(s,a)\sim d_{T}^{\mu}}\big{[}R^{*}(s,a)-\widehat{R}(s,a)\big{]}}\Bigg{|}.$ (13)
where $R^{*}(s,a)$ is the true reward model and coefficient $\mathscr{C}_{\mathcal{R}}(\pi)$ is upper bounded by $\|d_{T}^{\pi}/d_{T}^{\mu}\|_{\infty}$.
Next, similar to [20, 18], we use $\varepsilon$-bracketing number to measure the complexity of reward function class $\mathcal{R}$, which can be defined as
@@
Under Assumption 1, for any $\delta\in(0,1]$, the policy $\widehat{\pi}$ learned by LEASE, with high probability $1-\delta$, satisfies that $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$ is upper bound by
@@
where $C>0$ is the absolute constant defined in Eq. (2), $N$ is the size of preference dataset, and $L$ is the length of trajectory. The term $\xi$ is performance gap depending on offline algorithm itself.
The proof of Theorem 2 can be found in Appendix A-C. The $\xi$ is constant when offline algorithm is determined. Therefore, the tighter bound can be achieved through reducing performance gap caused by reward model (see Proposition 3). It can be reduced to small value $\varpi$ with sample complexity of
@@
$+\sqrt{\frac{R_{max}^{2}\log\big{(}1/\delta\big{)}}{L}}\Bigg{)}^{2}\Bigg{)},$
where concentrability coefficient $\mathscr{C}_{\mathcal{R}}(\widehat{\pi})$ can become smaller when the learned policy $\widehat{\pi}$ is close to behavior policy $\mu$. Bracketing number $\log(\mathcal{N}_{\mathcal{R}}(1/N))$ measures the complexity of function class $\mathcal{R}$ and takes $\widetilde{\mathcal{O}}(d)$ in linear reward model [18]. Notably, LEASE can learn accurate reward model through data augmentation under fewer $N_{l}$ preference data ($N_{l}<N$). Thus, the performance gap can become tighter under fewer data, that is LEASE can have higher sample efficiency compared with [18]. More discussion can be found in Section VIII-A.
VII Experiments
This section gives the details of experiments and focuses on answering the following questions:
  • $\mathbf{Q_{1}}$: How well the accuracy of reward model through data augmentation under fewer preference data?
  • $\mathbf{Q_{2}}$: How well does LEASE perform compared with offline PbRL baseline in standard benchmark tasks?
  • $\mathbf{Q_{3}}$: How the proposed selecting mechanism and parameter settings affect the performance of reward and policy?
We answer these questions on two distinct domains (mujoco and adroit) within the D4RL benchmark [22], covering several locomotion and manipulation tasks. The code for LEASE is available at anonymous.link.
VII-A Experiment Setup
@@
Experimental Datasets. The offline dataset and preference dataset of experiments originate from [22] and [21], respectively. The offline datasets are based on the D4RL dataset [22]. We select medium and medium-expert two types dataset for Mujoco tasks, and human and expert for Adroit tasks. The difference between different dataset in certain task lies in the collected policy. Fig. 2 depicts the six tasks for D4RL benchmark. For three mujoco tasks, the goal is to control the robot’s various joints to achieve faster and more stable locomotion. For adroit tasks, they involve controlling a 24-DoF simulated Shadow Hand robot on a sparse reward, high-dimensional robotic manipulation task.
The preference dataset are from the previous work [21], where they provided baseline for offline PbRL based on $2000$ preference data labeled by two types (We denote this method as URLHF). One is crowd-sourced labels obtained by crowd-sourcing, and the other is synthetic labels based on ground truth reward. To test LEASE performance for two types preference data, the labels of preference is from human-realistic feedback for mujoco tasks, and ground truth reward for adroit tasks. This paper aims to test the performance under fewer preference data without online interaction. We take the first $100$ data of [21] as $2000$ preference dataset, and denote method without data augmentation as FEWER.

### v2

@@ IV-A Uncertainty-aware Pseudo-label Selection
where $\kappa_{p}$ and $\kappa_{\tau}$ are thresholds of confidence and uncertainty respectively. $f(\sigma_{0}^{u},\sigma_{1}^{u})=1$ denotes the generated data $(\sigma_{0}^{u},\sigma_{1}^{u})$ is selected, and $f(\sigma_{0}^{u},\sigma_{1}^{u})=0$ denotes the data is not selected.
Intuitively, more rigorous data selection criteria, characterized by higher confidence and lower variance, lead to a reduction in the number of qualified samples. While such a strategy improves the quality of the generated data, it could lead to a narrower data distribution, which might adversely affect the precision of the trained reward model. Therefore, appropriate selection of parameters $\kappa_{p}$ and $\kappa_{\tau}$ is required to enhance the performance of the reward model. Then, combining Eq. (6) and the designed selection mechanism, the reward model can be optimized by minimizing
$\widehat{\mathcal{L}}_{R}(\psi)=\frac{1}{\widetilde{N}_{u}}\sum\nolimits_{j=1}^{{N}_{u}}f\left(\sigma_{0}^{u},\sigma_{1}^{u}\right)^{(j)}L\left((\sigma_{0}^{u},\sigma_{1}^{u})^{(j)},\widehat{y}^{j}\right)$ (9)
@@
where $\widetilde{N}_{u}$ is the number of generated dataset after screening, and $\widehat{\mathcal{L}}_{R}(\psi)$ is the empirical risk. It contains two parts: labeled loss $\widehat{\mathcal{L}}_{l}(\psi)$ and unlabeled loss $\widehat{\mathcal{L}}_{u}(\psi)$. Note that the label in the unlabeled loss is a pseudo-label, which may be different from the true label. Therefore, there is a gap between the unlabeled loss with pseudo-label $\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$. Before bounding this gap, we first give the following assumption without loss of generality.
@@
Note that the bound of pseudo-label error rate $\eta$ of the generated data filtered by the proposed selection mechanism is significantly lower than that of the unfiltered data, as the mechanism effectively eliminates low-confidence and high-uncertainty samples. This observation is experimentally validated in Table II and Fig. 3. Then, based on the above assumption, we bound the difference between the unlabeled loss with pseudo-label $\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ and give the following proposition.
@@
$\big|\widehat{\mathcal{L}}_{u}(\psi)-\widehat{\mathcal{L^{\prime}}}_{u}(\psi)\big|\leq\eta\Omega.$ (10)
[…]
IV-B Generalization Bound for Reward Model
This part studies the generalization performance of the reward model. Before developing generalization bound, we define the expected error $\mathcal{L}_{R}(\psi)$ with respect to the reward model $R(s,a;\psi)$: $\mathcal{L}_{R}(\psi)=\mathbb{E}_{(\sigma_{0},\sigma_{1},y)\sim\mathcal{D}}[L((\sigma_{0},\sigma_{1}),y)]$, where $\mathcal{D}$ is data distribution. The reward model is trained through minimizing the empirical error $\widehat{\mathcal{L}}_{R}(\psi)$ in Eq. (9). Through developing generalization error bound, we can analyze the factors that influence generalization ability.
Similar to the previous generalization bound works [29], Rademacher complexity, which measures the richness of a certain hypothesis space [19], is introduced first. The definition is given below:
@@ IV-B Generalization Bound for Reward Model
$\widehat{\Re}_{\widehat{\mathcal{S}}}(\mathcal{F})=\mathbb{E}_{\bm{\sigma}}\big[\sup_{f\in\mathcal{F}}\frac{1}{N_{s}}\sum\nolimits_{i=1}^{N_{s}}\sigma_{i}f\left(z_{i}\right)\big],$ (11)
@@
In general, lower Rademacher complexity corresponds to better generalization performance. Then, based on the above definition and Proposition 1, we develop the new generalization bound for the reward model trained through Eq. (9). The specific theorem is given below.
@@
The proof of Theorem 1 can be found in Appendix B. Theorem 1 indicates that the expected error $\mathcal{L}_{R}(\psi)$ is bounded by empirical error $\widehat{\mathcal{L}}_{R}(\psi)$, pseudo-labeling error $\eta\Omega$, Rademacher complexity and constant terms. The constant term can become small when the number of unlabeled data increases, but it may cause unstable training when labels are inaccurate. According to Proposition 1, empirical error and pseudo-labeling error can be reduced through the selection mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$, thus, a better generalization ability, that is tighter upper bound, can be achieved. Moreover, Theorem 1 has generality and can apply to methods that train a reward model using pseudo-labels.
V Policy Learning
To solve the problem of high cost for labeling preference, we propose a offLine prEference-bAsed RL with high Sample Efficiency algorithm (LEASE). This section aims to describe the details of LEASE and establish the theory for a safe policy improvement guarantee. In this paper, we directly use the proposed offline RL algorithms, such as CQL [4] and IQL [25], to perform policy learning, and do not improve the offline RL algorithm itself.
V-A The Implementation Details of LEASE
Most offline RL algorithms are based on Eq. (1). They proposed various algorithms to solve the problem of distribution shift. In an offline PbRL setting, the reward is unknown to the agent, and the accuracy of the reward model directly influences the performance of the policy. Moreover, the unlabeled data are generated through the current policy interaction with the learned transition model. The learned policy affects the distribution of the generated unlabeled preference dataset. In practical implementation, the policy, $Q$-function, reward and transition model are parameterized by related neural networks $\pi_{\theta}$, $Q_{\omega}$, $R_{\psi}$ and $T_{\varphi}$, respectively.
Algorithm 1 gives pseudo-code for LEASE. The goal of LEASE is to learn a better reward model $R_{\psi}$ from a smaller labeled preference dataset $\mathcal{D}_{l}$ and learn a better policy $\pi_{\theta}$ from an offline dataset $\mathcal{D}_{\text{offline}}$ without interaction with the environment. The key to LEASE is how to train the reward model in the policy learning process. The proposed algorithmic framework comprises three key components: reward model training, transition model training, and policy training. Initially, the transition model is trained using limited offline datasets, while the reward model undergoes pretraining with a small-scale preference dataset. The reward update condition is triggered when the number of unlabeled data in the buffer reaches its maximum capacity. Subsequently, the policy training phase commences. During this phase, the transition model is periodically employed to generate unlabeled synthetic data, and the reward model undergoes iterative updates to enhance its performance.
V-B Safe Policy Improvement Guarantee
Offline RL aims to guarantee $\xi$-safe policy improvement over the behavior policy $\mu$ (the policy used to collect offline dataset), i.e. $J(\widehat{\pi},R^{*})\geq J(\mu,R^{*})-\xi$ [4]. For offline PbRL, we further investigate whether similar policy improvement guarantees can be established. Therefore, this part aims to develop a theoretical guarantee of policy improvement for LEASE, that is, giving the bound for $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$. Then, based on the bounds of policy improvement, we further analyze the sample complexity for LEASE when the reward model is unknown.
@@ V-B Safe Policy Improvement Guarantee
5:  if $t$ % rollout frequency =0=0 then
6:   Generate data by transition model ${T}_{\varphi}$, and add transition trajectory into $\mathcal{D}_{u}$.
@@
8:  if $t$ % reward update frequency =0=0 then
9:   Generate pseudo label for unlabeled dataset $\mathcal{D}_{u}$, and screen unlabeled data by Eq. (8).
10:   Train the reward model based on $\mathcal{D}_{l}$ and $\mathcal{D}_{u}$ according to Eq. (9).
11:  end if
12:  Train policy using offline RL algorithms.
13:end for
@@
Before deriving the bounds of policy improvement, we first define a new single-policy concentrability coefficient for PbRL based on the state-action pair. This coefficient can quantify the discrepancy between the offline distribution and the visitation distribution of policy $\pi$.
@@
$\mathscr{C}_{\mathcal{R}}(\pi)=\Bigg|\frac{\mathbb{E}_{(s,a)\sim d_{T}^{\pi}}\big[R^{*}(s,a)-\widehat{R}(s,a)\big]}{\mathbb{E}_{(s,a)\sim d_{T}^{\mu}}\big[R^{*}(s,a)-\widehat{R}(s,a)\big]}\Bigg|,$ (13)
where $R^{*}(s,a)$ is the true reward model and $\hat{R}(s,a)$ is the learned reward model.
[…]
@@
Bracketing number $\mathcal{N}_{\mathcal{R}}(\varepsilon)$ measures the complexity of function class $\mathcal{R}$ and takes $\widetilde{\mathcal{O}}(d)$ in linear reward model [18]. Then, before giving the theorem of policy improvement, we first assume the reward class $\mathcal{R}$ is bounded, that is
Assumption 3 (Boundedness).
@@
Under Assumptions 1 and 3, for any $\delta\in(0,1]$, the policy $\widehat{\pi}$ learned by LEASE, with high probability $1-\delta$, satisfies that $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$ is upper bound by
@@
where $C>0$ is the absolute constant defined in Lemma 2, $N$ is the size of preference dataset, and $L$ is the length of trajectory. The term $\xi$ is the performance gap depending on the offline algorithm itself.
The proof of Theorem 2 can be found in Appendix C. The bounds of the PbRL policy learned by LEASE comprise two distinct components: the first term $\xi$ stems from the inherent gap of the offline RL itself, and the second arises from the discrepancy induced by the reward model. The first term $\xi$ is constant when the offline algorithm is determined. The second term is determined by the performance of the reward model. Therefore, the tighter bound can be achieved through reducing the performance gap caused by the reward model (see Proposition 3). It can be reduced to a small value $\varpi$ with a sample complexity of
@@
$+\sqrt{\frac{R_{max}^{2}\log\big(1/\delta\big)}{L}}\Bigg)^{2}\Bigg).$
[…]
VI Experiments
This section gives the details of experiments and presents comprehensive experimental evaluations of the proposed algorithm, including the accuracy assessment of reward model learning, performance analysis of policy learning, and detailed parameter sensitivity studies. We focus on answering the following questions:
  • $\mathbf{Q_{1}}$: How well does the accuracy of the reward model through data augmentation under fewer preference data?
  • $\mathbf{Q_{2}}$: How well does LEASE perform compared with the offline PbRL baseline in standard benchmark tasks?
  • $\mathbf{Q_{3}}$: How does the proposed selection mechanism and parameter settings affect the performance of reward and policy?
We answer these questions on two distinct domains (mujoco and adroit), covering several locomotion and manipulation tasks, within the D4RL benchmark [22]. The code for LEASE is available at Code.
VI-A Experiment Setup
@@
VI-A1 Experimental Datasets
The offline dataset and preference dataset of experiments originate from [22] and [21], respectively. The offline datasets are based on the D4RL dataset [22]. We select medium and medium-expert two types of datasets for Mujoco tasks and human and expert for Adroit tasks. The difference between datasets in a certain task lies in the policies they collect. The Mujoco tasks include halfcheetah-v2, hopper-v2, and walker2d-v2. The Adroit tasks include pen-v1, door-v1, and hammer-v1. Fig. 2 depicts the six tasks for the D4RL benchmark. For three mujoco tasks, the goal is to control the robot’s various joints to achieve faster and more stable locomotion. For adroit tasks, they involve controlling a 24-DoF simulated Shadow Hand robot on a sparse reward, high-dimensional robotic manipulation task.
The preference dataset is from the previous work [21], where they provided the baseline for offline PbRL based on $2000$ preference data labeled by two types (We denote this method as URLHF). One is crowd-sourced labels obtained by crowd-sourcing, and the other is synthetic labels based on ground truth reward. To test LEASE performance for two types of preference data, the labels of preference are from human-realistic feedback for MuJoco tasks, and ground truth reward for Adroit tasks. This paper aims to test the performance under fewer preference data without online interaction. We take the first $100$ data of [21] as $2000$ preference dataset, and denote the method without data augmentation as FEWER.
VI-A2 Experimental Details
The Offline PbRL setting involves two steps: reward learning and policy learning. Here, we directly use offline RL algorithms CQL [4] and IQL [25] to perform policy learning. The performance of these algorithms is evaluated based on the cumulative return. For comparison, the scores are normalized between 0 (random policy score) and 100 (expert policy score) [22]. The normalized score $\tilde{J}$ is computed by:
$\tilde{J}=\frac{J-J_{r}}{J_{e}-J_{r}}\times 100,$ (16)
where $J_{r}$, $J_{e}$, and $J$ are the expected returns of a random policy, an expert policy, and the trained policy by offline RL algorithms, respectively. Table I gives the hyperparameter configuration for reward and transition training, and policy learning. The details are given below: